OpenStudy (anonymous):

Show all work. I don't get how to do these. Thanks in advance! :) Will give medal. 1. Multiply and simplify. a) (b + 2)(b + 2) b) (x + 3)(x – 3) 2. Simplify. a) (3xy3)(2x3y) b) 2y(2y + 3y2) – y2(3 – y)

OpenStudy (texaschic101):

1. FOIL method First - outside - inside- last (b + 2)(b + 2) first -- multiply first terms in each set...b * b = b^2 outside -- multiply outside terms in each set ... b * 2 = 2b inside -- multiply inside terms in each set .. 2 * b = 2b last -- multiply last terms in each set .. 2 * 2 = 4 now put it all together b^2 + 2b + 2b + 4 -- combine like terms b^2 + 4b + 4 <== your answer you would do the other one on part 1 the same way. ======================= not too good with part 2...sorry

OpenStudy (anonymous):

Thank you.